40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 33     Entry time: Fri Jan 15 11:35:17 2010
Author: Frank 
Type: Laser 
Category: RefCav 
Subject: refcav update 

the last couple of days we fixed a couple of thinks:

We found a wrong calibration of the temp readout of the first refcav, which gave us a wrong residual noise level compared to the other one. The absolute value was almost correct so we didn't realize this before. Now we are limited by the DAQ noise so the next step will be improving the gain settings in the temp sensor readout-box.

A big problem is the changing room temperature as soon someone opens the door even for only a couple of seconds. The temp control of the lab goes crazy and changes a couple of degrees and is oscillating half a day after that. The delay to the cavities is about 30min and a couple of deg changes of the room temp also change the cavity temperature, only a couple of tens of mK but enough to shift both cavities away from each other as the room temp couples different into both due to the different thermal insulation (insulation and time constants are different). We opened the room temp sensor and adjusted the temperature to i higher level, removed surrounding parts to get a better air flow to the sensor. We don't understand why opening the door has such a large effect on the room temp as 10 seconds of open door don't change the room temp by 2 degC or so. So if anybody has an idea plz let us know even if it seems to be stupid.

In order to improve the temp stability of the analyzer cavity we added a second layer of insulation and wrapped the whole thing in aluminum foil, see picture below. The insulation is 2" thick, except a small part at the large flanges at the end where its only 1" thick. We also changed the gain of the temp sensor readout in order to reduce the influence of DAQ noise. If nobody is working in the lab we have a stability of about 6mKpp within hours. As soon someone is working in one of the labs this changes to tens of mK for working in other labs to about a hundred mK if working in the PSL lab. We are currently working on improved servo settings...

100_0542.JPG

ELOG V3.1.3-